OpenStudy (anonymous):

Peter's walking speed is 1.5 m/s, while Davis's walking speed is 1 m/s. (a) If they start walking away from each other, how long will it be for them to be 100m apart? (b) If they start walking in the same direction at the same time, how long will it be for them to be 100 m apart?

OpenStudy (anonymous):

A) Let the time taken for them to be 100 m apart = x seconds in this time Peter covers 1.5x m and Davis covers 1x m ATQ, 1.5x + x = 100 2.5x = 100 x = 100/2.5 = 40 so they will be 100 m apart after 40 seconds

OpenStudy (anonymous):

A) When they walk in the same direction, Let the time taken for them to be 100 m apart = x seconds in this time Peter covers 1.5x m and Davis covers 1x m Since Peter's speed is more, he will be ahead ATQ, 1.5x - x = 100 0.5x = 100 x = 100/0.5 = 200 so if they walk in the same direction, they will be 100 m apart after 200 seconds i.e. 3 min 20 seconds